OpenStudy (anonymous):

Graphing inequalities. 2x + 4y is equal or greater than 4 4x + y is equal or greater than 8 Help me please?

OpenStudy (anonymous):

Can someone explain this to me step by step? I do not understand.

jimthompson5910 (jim_thompson5910):

First solve each for y 2x + 4y >= 4 4y >= - 2x + 4 y >= - 2x/4 + 4/4 y >= -1/2x + 1

jimthompson5910 (jim_thompson5910):

4x + y >= 8 y >= -4x + 8

jimthompson5910 (jim_thompson5910):

so 2x + 4y >= 4 4x + y >= 8 turns into y >= -1/2x + 1 y >= -4x + 8

jimthompson5910 (jim_thompson5910):

To graph y >= -1/2x + 1, you follow these steps 1) graph the line y = -1/2x + 1 2) make this line a solid line 3) shade above the line to graph y >= -1/2x + 1

jimthompson5910 (jim_thompson5910):

To graph y >= -4x + 8, you follow these steps 1) graph the line y = -4x + 8 2) make this line a solid line 3) shade above the line to graph y >= -4x + 8

jimthompson5910 (jim_thompson5910):

does that help?

OpenStudy (anonymous):

Where do I plot the points?

jimthompson5910 (jim_thompson5910):

graphing the line y = -1/2x + 1 the y - intercept is (0,1) the slope is -1/2 so you start at (0,1) and go down 1 and to the right 2 units to get to the next point (2,0) then draw a line through (0,1) and (2,0) to graph y = -1/2x + 1

jimthompson5910 (jim_thompson5910):

do the same for the next line as well

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

yw